A particular solution of the nonhomogeneous differential equation y00 − 2 y0 − 15 y = 2 cos 3x + 5 e5x + 2 will have the form:
Answer #1

you_24--159 = 26053% + 50544-2 Yp = ? (0²2D-15)y Up = 2.2053x+5652+2 200S3X +5052+2 D² 20-15 - Lesx + 2 3-2 D-15 2 5. COS + JHere D.-20-15= D 5D+3D -15 • DCD-5)+3 CD-5)=(1-5) (D+3) = 57 est = es D2D-15 be D+3 013) (0-5) xest xet = D+3 =Ine 54 13 5+3